Please help
Three rectangular prisms have a combined volume of 518 cubic feet. Prism A has one-third the volume of Prism B, Prism B and C have equal volume. What is the volume of each prism?

Answer by mananth(16949)

Three rectangular prisms have a combined volume of 518 cubic feet. Prism A has one-third the volume of Prism B, Prism B and C have equal volume. What is the volume of each prism
Let B & C have volumes = x
prism A will have x/3
x/3 + x + x = 518
7x/3 = 518
7x = 518 *3
x=74*3
x= 222 cu.ft ( B & C)
Vol A = 74 cu.ft